a rectangular piece of cardboard is 4 inches longer than it is wide. From each of its corners, a square piece 2 inches on a side is cut out. The flaps are then turned up to form an open box, which has a volume of 42 cubic inches. Find the length and width of the original piece of cardboard.
Answer by stanbon(75887)

a rectangular piece of cardboard is 4 inches longer than it is wide. From each of its corners, a square piece 2 inches on a side is cut out. The flaps are then turned up to form an open box, which has a volume of 42 cubic inches. Find the length and width of the original piece of cardboard.
------------------
Width = x
Length = x+4
Volume = 2(x-4)(x+4-4) = 2(x-4)(x) = 42 in^3
-------
x^2 -4x- 21 = 0
Factor:
(x-7)(x+3) = 0
x = 7 inches
----
width = x = 7
length = x+4 = 11
========================
Cheers,
Stan H.
========================